OpenStudy (anonymous):

Can some one help me, I'm trying to find the percentage increase between 0.69 $1.55 $2.60 $4.23 $5.39

OpenStudy (dariusx):

0.69x = 1.55 x = 1.55/0.69 = 2.24637 now converting the number into percents x = 2.24637 * 100 = 224.64 % that's a 224.64% increase from number 0.69 to 1.55 try to do others yourself